OpenStudy (anonymous):

i am currently in college algebra and im doing inverse functions f(x) and g(x) f(x) 4x-16, g(x) x/4 +4 and im having problems with simplifying 4x-4/4 +4 can you please help

OpenStudy (anonymous):

i have to substitue the 4x-16/ 4 in for gx

OpenStudy (zzr0ck3r):

f(x) = 4x-16 lets call this y = 4x - 16 now we want to solve for x in terms of y add 16 to both sides y+16 = 4x now devide away the 4 (y/4)+4 = x so now "switch the x and y" (x/4) + 4 = y f(x) inverse is f^(-1)(x) = (x/4) + 4

OpenStudy (anonymous):

ok i think i understand but im doing one to one functions to see if they inverses of each other

OpenStudy (anonymous):

pluging one function in for the other